OpenStudy (anonymous):

limit x-> 1 x^3-1 / x^3 - 6x^2 + 11x -6

OpenStudy (across):

\[\lim_{x\to1}\frac{x^3-1}{x^3}-6x^2+11x-6\]is the same thing as\[\lim_{x\to1}\frac{x^3-1}{x^3}-\lim_{x\to1}6x^2+\lim_{x\to1}11x-\lim_{x\to1}6.\]

OpenStudy (anonymous):

so ? i don't understand.

OpenStudy (across):

Substitute 1's into the x's and solve the equation...

OpenStudy (anonymous):

it will be 0/0 = undetermined

OpenStudy (anonymous):

(x^3-6x^2+11x-6)=(x-1)(x-2)(x-3) lim x->1(x^3-1)/(x^3-6x^2+!1x-6)=limx->1(x^2+x+1)/(x-2)(x-3)=3/2